OpenStudy (anonymous):

abase ballisthrown vertivaly upward with velocty20m/s.find 1)maximuheigh timetakatoreachmaximum

OpenStudy (anonymous):

h=(v^2) / 2g h=20m use v2=u2-2gh [u=0] then calculate h

OpenStudy (anonymous):

t=sqrt (2h/g) t=4secs